60 people can finish a work in 15 days. After doing 6 days of work, 12 people added to the group for completing the same work. In how many days, they will complete the work?
Found 2 solutions by ankor@dixie-net.com, richwmiller:
Answer by ankor@dixie-net.com(22740)

60 people can finish a work in 15 days
Find the no. of man-days required to complete the work
60 * 15 = 900 man-days
:
After doing 6 days of work, 12 people added to the group for completing the same work.
In how many days, they will complete the work?
Let x = no. days to complete the job after 12 more were added
60(x+6) + 12x = 900
60x + 360 + 12x = 900
60x + 12x = 900 - 360
72x = 540
x = 540/72
x = 7.5 additional days:
:
Confirm this
Original 60 worked 6 + 7.5 = 13.5 days
60 * 13.5 = 810 man-days
12 * 7.5 = 90 man-days
------------------------
total man-days: 900

Answer by richwmiller(17219)

man-days 60*15=900
man days worked=60*6=360
man days needed=900-360=540
new number of workers=60+12=72
man days needed/new number of workers 540/72= 7.5 days needed